# I Predict Therefore I Am: Is Next Token Prediction Enough to Learn Human-Interpretable Concepts from Data?

### v2

@@ Appendix B Limitations
    $h_{y_{i}}-h_{y_{0}}\xrightarrow[\epsilon\to 0]{}0.$
  • Recall the definition of the quantity $\mathbf{b}_{y}$. For any pair of labels $y_{i}$ and $y_{0}$, the difference between components is:
    $b_{y_{i}}-b_{y_{0}}=\mathbb{E}_{p(\mathbf{c}\mid y=y_{i})}\left[\log p(y=y_{i}\mid\mathbf{c})\right]-\mathbb{E}_{p(\mathbf{c}\mid y=y_{0})}\left[\log p(y=y_{0}\mid\mathbf{c})\right].$ (34)
    Since the generative process $\mathbf{g}:\mathbf{c}\rightarrow(\mathbf{x},y)$ is approximately invertible, i.e., $\epsilon\to 0$, the conditional $p(y\mid\mathbf{c})$ becomes sharply peaked. In particular, there exists a dominant latent mode $\mathbf{c}^{*}_{y_{i}}$ such that
    $p(y=y_{i}\mid\mathbf{c}^{*}_{y_{i}})=1-\epsilon^{\prime},\quad p(y\neq y_{i}\mid\mathbf{c}^{*}_{y_{i}})=\epsilon^{\prime},$
    where $\epsilon^{\prime}\to 0$ as $\epsilon\to 0$.
    Therefore, the expected log-probability becomes:
    $\mathbb{E}_{p(\mathbf{c}\mid y=y_{i})}\left[\log p(y=y_{i}\mid\mathbf{c})\right]$ $=p(\mathbf{c}^{*}_{y_{i}}\mid y=y_{i})\log p(y=y_{i}\mid\mathbf{c}^{*}_{y_{i}})$
    $\quad+\sum_{\mathbf{c}^{\prime}\neq\mathbf{c}^{*}_{y_{i}}}p(\mathbf{c}^{\prime}\mid y=y_{i})\log p(y=y_{i}\mid\mathbf{c}^{\prime})$ (35)
    $\approx\log(1-\epsilon^{\prime})+\sum_{\mathbf{c}^{\prime}\neq\mathbf{c}^{*}_{y_{i}}}p(\mathbf{c}^{\prime}\mid y=y_{i})\log\epsilon^{\prime}.$ (36)
    As $\epsilon^{\prime}\to 0$, we have $\log(1-\epsilon^{\prime})\to 0$. Further, since $p(\mathbf{c}^{\prime}\mid y=y_{i})$ is bounded, the second tends to be $\log\epsilon^{\prime}$, which only depends on $\epsilon^{\prime}$. Therefore, the difference becomes:
    $b_{y_{i}}-b_{y_{0}}\xrightarrow[\epsilon\to 0]{}0.$
@@
According to the assumption in Theorem 3.1, there exists a matrix $\mathbf{A}$ such that:
$\mathbf{f}_{x}(\mathbf{x})\approx\mathbf{A}[\log p(\mathbf{c}=\mathbf{c}_{i}|\mathbf{x})]_{\mathbf{c}_{i}}$ (37)
where $[\cdot]_{\mathbf{c}_{i}}$ denotes concatenating $\log p(\mathbf{c}_{i}|\mathbf{x})$ over all possible values of $\mathbf{c}$.
@@
Since $1-p(\mathbf{c}=\mathbf{c}^{*}_{\mathbf{x}}|\mathbf{x})=\epsilon_{\mathbf{x}}$, when $\epsilon_{\mathbf{x}}\rightarrow 0$, we have:
@@
$\left[\log p(\mathbf{c}={\mathbf{c}_{i}}|\mathbf{x})\right]_{\mathbf{c}_{i}}\approx[\log o(\epsilon_{\mathbf{x}}),...,\log(1-\epsilon_{\mathbf{x}}),...,\log o(\epsilon_{\mathbf{x}})],$ (38)
where where $\log(1-\epsilon_{x})$ appears at the index corresponding to $\mathbf{c}_{x}^{*}$, and other entries are $\log o(\epsilon_{x})$.
Now consider two inputs $\mathbf{x}_{0}$ and $\mathbf{x}_{1}$, with their most likely concepts being $\mathbf{c}_{x_{0}}^{*}$ and $\mathbf{c}_{x_{1}}^{*}$, corresponding to indices $k_{0}$ and $k_{1}$, respectively, with $k_{0}\neq k_{1}$. We have:
$f_{\mathbf{x}}(\mathbf{x}_{1})-f_{\mathbf{x}}(\mathbf{x}_{0})$ $\approx\mathbf{A}\left(\left[\log p(\mathbf{c}=\mathbf{c}_{j}\mid\mathbf{x}_{1})\right]_{\mathbf{c}_{j}}-\left[\log p(\mathbf{c}=\mathbf{c}_{j}\mid\mathbf{x}_{0})\right]_{\mathbf{c}_{j}}\right)$ (39)
$\approx\mathbf{A}\left(\left[0,\ldots,\underbrace{\log(1-\epsilon_{\mathbf{x}_{1}})-\log o(\epsilon_{\mathbf{x}_{0}})}_{k_{1}},\ldots,\underbrace{\log o(\epsilon_{\mathbf{x}_{1}})-\log(1-\epsilon_{\mathbf{x}_{0}})}_{k_{0}},\ldots,0\right]\right)$ (40)
Only the $k_{1}$-th and $k_{0}$-th positions are nonzero, so this can be written as:
$f_{\mathbf{x}}(\mathbf{x}_{1})-f_{\mathbf{x}}(\mathbf{x}_{0})\approx\left(\underbrace{\log(1-\epsilon_{\mathbf{x}_{1}})-\log o(\epsilon_{\mathbf{x}_{0}})}_{s_{1}}\right)\mathbf{A}_{:,k_{1}}+\left(\underbrace{\log o(\epsilon_{\mathbf{x}_{1}})-\log(1-\epsilon_{\mathbf{x}_{0}})}_{s_{2}}\right)\mathbf{A}_{:,k_{0}}$
Then we obtain:
$f_{\mathbf{x}}(\mathbf{x}_{1})-f_{\mathbf{x}}(\mathbf{x}_{0})\approx s_{1}\mathbf{A}_{:,k_{1}}-s_{2}\mathbf{A}_{:,k_{0}}$
Since the inputs $\mathbf{x}_{1}$ and $\mathbf{x}_{0}$ differ only in the $i$-th concept variable $c_{i}$, and share identical values for all other concept variables $\mathbf{c}_{-i}$. Specifically, let:
  • $\mathbf{x}_{1}$ corresponds to concept assignment $\mathbf{c}_{1}^{*}=(c_{i}=1,\mathbf{c}_{-i}^{*})$;
  • $\mathbf{x}_{0}$ corresponds to concept assignment $\mathbf{c}_{0}^{*}=(c_{i}=0,\mathbf{c}_{-i}^{*})$;
  • The conditional probabilities satisfy $p(\mathbf{c}_{1}^{*}\mid\mathbf{x}_{1})=1-\epsilon_{\mathbf{x}_{1}}$ and $p(\mathbf{c}_{0}^{*}\mid\mathbf{x}_{0})=1-\epsilon_{\mathbf{x}_{0}}$, where $\epsilon_{\mathbf{x}_{1}},\epsilon_{\mathbf{x}_{0}}\to 0$.
Now consider the conditional probability $p(c_{i}=1\mid\mathbf{x}_{1})$:
$p(c_{i}=1\mid\mathbf{x}_{1})=\sum_{\mathbf{c}_{-i}}p(c_{i}=1,\mathbf{c}_{-i}\mid\mathbf{x}_{1})=p(\mathbf{c}_{1}^{*}\mid\mathbf{x}_{1})+\sum_{\mathbf{c}_{-i}^{\prime}\neq\mathbf{c}_{-i}^{*}}p(c_{i}=1,\mathbf{c}_{-i}^{\prime}\mid\mathbf{x}_{1}).$
Because $p(\mathbf{c}_{1}^{*}\mid\mathbf{x}_{1})=1-\epsilon_{\mathbf{x}_{1}}$ and the remaining probability mass is small (order $\epsilon_{\mathbf{x}_{1}}$), we obtain:
$p(c_{i}=1\mid\mathbf{x}_{1})\approx 1-\epsilon_{\mathbf{x}_{1}}.$
Thus:
$s_{1}=\log(1-\epsilon_{\mathbf{x}_{1}})-\log o(\epsilon_{\mathbf{x}_{0}})\approx\log(1-\epsilon_{\mathbf{x}_{1}})\approx\log p(c_{i}=1\mid\mathbf{x}_{1}),$
and similarly,
$s_{2}\approx\log p(c_{i}=0\mid\mathbf{x}_{0}).$
@@
Let $\mathbf{f}_{\mathbf{x}}(\mathbf{x})\approx\mathbf{A}[\log{p(\mathbf{c}=\mathbf{c}_{i}|\mathbf{x})}]_{\mathbf{c}_{i}}$ as given in Theorem 3.1. For a data pair $(\mathbf{c},\mathbf{x})$, where $\mathbf{x}$ is generated by the latent variable $\mathbf{c}$, the representations $\mathbf{f}_{\mathbf{x}}(\mathbf{x})$ are passed to a linear classifier with weights $\mathbf{W}$.
@@
$\mathbf{z}\approx\mathbf{W}\big{(}\mathbf{A}[\log{p(\mathbf{c}=\mathbf{c}_{i}|\mathbf{x})}]_{\mathbf{c}_{i}}\big{)}.$ (41)
For correct classification under cross entropy loss, the logits must match the true log-probabilities:
$\mathbf{z}=\log{p(\mathbf{c}|\mathbf{x})}.$ (42)
@@
$\mathbf{W}\mathbf{A}\approx\mathbf{I},$ (43)

### v3

@@ Appendix B Limitations
$h_{y_{i}}-h_{y_{0}}\xrightarrow[\epsilon\to 0]{}0.$ (33)
@@
Proof.
We first prove that: when $\epsilon_{\mathbf{x}}\to 0$, i.e., when $p(\mathbf{c}\mid\mathbf{x})$ becomes sharply peaked at $\mathbf{c}_{\mathbf{x}}^{*}$, we can approximate:
$p(\mathbf{c})\approx p(c^{i}\mid\mathbf{x})\cdot p(\mathbf{c}^{-i}\mid\mathbf{x}),$ (34)
where $\mathbf{c}^{-i}$ denotes all concepts except $c^{i}$. To this end, we analysis their KL(Kullback–Leibler) residual term:
$\text{Residual}:=D_{\text{KL}}\left(p(\mathbf{c}\mid\mathbf{x})\;\big{\|}\;p(c^{i}\mid\mathbf{x})p(\mathbf{c}^{-i}\mid\mathbf{x})\right).$ (35)
@@
Since when $\epsilon_{\mathbf{x}}\to 0$, $p(\mathbf{c}\mid\mathbf{x})$ is sharply peaked at a particular configuration $\mathbf{c}_{x}^{*}=(c^{i*},\mathbf{c}^{-i*})$, i.e.:
$p(\mathbf{c}\mid\mathbf{x})=\begin{cases}1-\epsilon_{\mathbf{x}},&\text{if }\mathbf{c}=\mathbf{c}_{x}^{*},\\ \epsilon_{\mathbf{x}}\cdot r(\mathbf{c}),&\text{otherwise},\end{cases}$ (36)
where $\sum_{\mathbf{c}\neq\mathbf{c}_{x}^{*}}r(\mathbf{c})=1$.
Main Term in KL
The KL divergence is defined as:
$D_{\text{KL}}(p(\mathbf{c}\mid\mathbf{x})\|{p(c^{i}\mid\mathbf{x})p(\mathbf{c}^{-i}\mid\mathbf{x})})=\sum_{\mathbf{c}}p(\mathbf{c}\mid\mathbf{x})\log\frac{p(\mathbf{c}\mid\mathbf{x})}{p(c^{i}\mid\mathbf{x})p(\mathbf{c}^{-i}\mid\mathbf{x})}.$ (37)
The main contribution comes from $\mathbf{c}=\mathbf{c}_{x}^{*}$:
$(1-\epsilon_{\mathbf{x}})\log\frac{1-\epsilon_{\mathbf{x}}}{p(c^{i*}\mid\mathbf{x})\cdot p(\mathbf{c}^{-i*}\mid\mathbf{x})}.$ (38)
Note that:
$p(c^{i*}\mid\mathbf{x})=\sum_{\mathbf{c}^{-i}}p(c^{i*},\mathbf{c}^{-i}\mid\mathbf{x})\geq p(c^{i*},\mathbf{c}^{-i*}\mid\mathbf{x})=1-\epsilon_{\mathbf{x}},$ (39)
and similarly:
$p(\mathbf{c}^{-i*}\mid\mathbf{x})\geq 1-\epsilon_{\mathbf{x}}.$ (40)
Hence:
$p(c^{i*}\mid\mathbf{x})\cdot p(\mathbf{c}^{-i*}\mid\mathbf{x})\geq(1-\epsilon_{\mathbf{x}})^{2},$ (41)
$\Rightarrow\quad\frac{1-\epsilon_{\mathbf{x}}}{p(c^{i*}\mid\mathbf{x})\cdot p(\mathbf{c}^{-i*}\mid\mathbf{x})}\leq\frac{1}{1-\epsilon_{\mathbf{x}}}.$ (42)
Thus, the main term becomes:
$(1-\epsilon_{\mathbf{x}})\log\left(\frac{1}{1-\epsilon_{\mathbf{x}}}\right)=-(1-\epsilon_{\mathbf{x}})\log(1-\epsilon_{\mathbf{x}}).$ (43)
Tail Term
For $\mathbf{c}\neq\mathbf{c}_{x}^{*}$, the contribution to the KL divergence is:
$\epsilon_{\mathbf{x}}r(\mathbf{c})\log\left(\frac{\epsilon_{\mathbf{x}}r(\mathbf{c})}{\epsilon_{\mathbf{x}}^{2}r(c^{i})r(\mathbf{c}^{-i})}\right)=\epsilon_{\mathbf{x}}r(\mathbf{c})\log\left(\frac{1}{\epsilon_{\mathbf{x}}}\cdot\frac{r(\mathbf{c})}{r(c^{i})r(\mathbf{c}^{-i})}\right).$ (44)
Summing over all $\mathbf{c}\neq\mathbf{c}_{x}^{*}$, we obtain:
$\epsilon_{\mathbf{x}}\sum_{\mathbf{c}\neq\mathbf{c}_{x}^{*}}r(\mathbf{c})\log\left(\frac{1}{\epsilon_{\mathbf{x}}}\cdot\frac{r(\mathbf{c})}{r(c^{i})r(\mathbf{c}^{-i})}\right)=\epsilon_{\mathbf{x}}\log\frac{1}{\epsilon_{\mathbf{x}}}+\epsilon_{\mathbf{x}}\sum_{\mathbf{c}}r(\mathbf{c})\log\left(\frac{r(\mathbf{c})}{r(c^{i})r(\mathbf{c}^{-i})}\right).$ (45)
Since the second term is bounded, we conclude:
Tail $\text{Tail Term}=o(\epsilon_{\mathbf{x}}\log\epsilon_{\mathbf{x}}).$ (46)
which vanishes faster than the main term as $\epsilon_{\mathbf{x}}\to 0$.
Final Bound
Combining both contributions in Main Term and Tail Term, we get:
$D_{\mathrm{KL}}\left(p(\mathbf{c}\mid\mathbf{x})\,\big{\|}\,p(c^{i}\mid\mathbf{x})\cdot p(\mathbf{c}^{-i}\mid\mathbf{x})\right)\leq-(1-\epsilon_{\mathbf{x}})\log(1-\epsilon_{\mathbf{x}})+o(\epsilon_{\mathbf{x}}\log\epsilon_{\mathbf{x}}).$ (47)
@@
Here when $\epsilon_{\mathbf{x}}\to 0$, $D_{\mathrm{KL}}\left(p(\mathbf{c}\mid\mathbf{x})\,\big{\|}\,p(c^{i}\mid\mathbf{x})\cdot p(\mathbf{c}^{-i}\mid\mathbf{x})\right)\to 0$.
Based on Eq. (34), $\left[\log p(\mathbf{c}\mid\mathbf{x})\right]_{\mathbf{c}_{i}}$ can be rewritten as:
$\left[\log p(\mathbf{c}\mid\mathbf{x})\right]_{\mathbf{c}_{i}}\approx\mathbf{B}^{i}\left[\log p(c^{i}\mid\mathbf{x})\right]_{c^{i}}+\mathbf{B}^{-i}\left[\log p(\mathbf{c}^{-i}\mid\mathbf{x})\right]_{\mathbf{c}^{-i}}.$ (48)
Here, $\mathbf{B}^{i}$ and $\mathbf{B}^{-i}$ are binary broadcasting matrices that expand the marginal log-probability vectors $\left[\log p(c^{i}\mid\mathbf{x})\right]_{{c}^{i}}$ and $\left[\log p(\mathbf{c}^{-i}\mid\mathbf{x})\right]_{\mathbf{c}^{-i}}$ to the full configuration space $[\log p(\mathbf{c}\mid\mathbf{x})]_{\mathbf{c}_{i}}$, respectively.
As a result, for pair $\mathbf{x}_{0}$ and $\mathbf{x}_{1}$ that differ only in the $i$-th concept variable $c^{i}$, their difference in representation space is:
$\left[\log p(\mathbf{c}\mid\mathbf{x}_{1})\right]_{\mathbf{c}_{i}}-\left[\log p(\mathbf{c}\mid\mathbf{x}_{0})\right]_{\mathbf{c}_{i}}$
$\approx\mathbf{B}^{i}\left[\log p(c^{i}\mid\mathbf{x}_{1})-\log p(c^{i}\mid\mathbf{x}_{0})\right]_{c^{i}}+\mathbf{B}^{-i}\left[\log p(\mathbf{c}^{-i}\mid\mathbf{x}_{1})-\log p(\mathbf{c}^{-i}\mid\mathbf{x}_{0})\right]_{\mathbf{c}^{-i}}.$ (49)
We now show that:
$p(\mathbf{c}^{-i}\mid\mathbf{x}_{1})\approx p(\mathbf{c}^{-i}\mid\mathbf{x}_{0}),$ (50)
so the term $\mathbf{B}^{-i}\left[\log p(\mathbf{c}^{-i}\mid\mathbf{x}_{1})-\log p(\mathbf{c}^{-i}\mid\mathbf{x}_{0})\right]_{\mathbf{c}^{-i}}$ in Eq. (49) vanishes.
Recall Eq. (36), we have
$p(\mathbf{c}\mid\mathbf{x})=\begin{cases}1-\epsilon_{\mathbf{x}},&\text{if }\mathbf{c}=\mathbf{c}_{x}^{*}=(c^{i*},\mathbf{c}^{-i*}),\\ \epsilon_{\mathbf{x}}\cdot r(\mathbf{c}),&\text{otherwise},\end{cases}$ (51)
Then, for $\mathbf{c}^{-i}$, we have:
$p(\mathbf{c}^{-i}\mid\mathbf{x})=\sum_{c^{i}}p(c^{i},\mathbf{c}^{-i}\mid\mathbf{x})=\begin{cases}1-\epsilon_{\mathbf{x}},&\text{if }(c^{i},\mathbf{c}^{-i})=\mathbf{c}_{x}^{*},\\ \sum_{c^{i}}\epsilon_{\mathbf{x}}\cdot r(c^{i},\mathbf{c}^{-i}),&\text{otherwise}.\end{cases}$ (52)
This yields:
$p(\mathbf{c}^{-i*}\mid\mathbf{x})$ $=1-\epsilon_{\mathbf{x}}+\epsilon_{\mathbf{x}}\sum_{c^{i}\neq c^{i*}}r(c^{i},\mathbf{c}^{-i*})=1-\epsilon_{\mathbf{x}}+o(\epsilon_{\mathbf{x}}),$ (53)
$p(\mathbf{c}^{-i}\mid\mathbf{x})$ $=\epsilon_{\mathbf{x}}\sum_{c^{i}}r(c^{i},\mathbf{c}^{-i})=o(\epsilon_{\mathbf{x}}),\quad\text{for }\mathbf{c}^{-i}\neq\mathbf{c}^{-i*}.$ (54)
whether $\mathbf{c}^{-i}=\mathbf{c}^{-i*}$ (the dominant configuration) or $\mathbf{c}^{-i}\neq\mathbf{c}^{-i*}$ (a non-dominant configuration), we have for all $\mathbf{c}^{-i}$,
$p(\mathbf{c}^{-i}\mid\mathbf{x}_{1})=p(\mathbf{c}^{-i}\mid\mathbf{x}_{0})+o(\epsilon_{\mathbf{x}}),$ (55)
which implies:
$\log p(\mathbf{c}^{-i}\mid\mathbf{x}_{1})-\log p(\mathbf{c}^{-i}\mid\mathbf{x}_{0})=\log\left(1+\frac{o(\epsilon_{\mathbf{x}})}{p(\mathbf{c}^{-i}\mid\mathbf{x}_{0})}\right)=o(\epsilon_{\mathbf{x}}).$ (56)
Consequently,
$\mathbf{B}^{-i}\left[\log p(\mathbf{c}^{-i}\mid\mathbf{x}_{1})-\log p(\mathbf{c}^{-i}\mid\mathbf{x}_{0})\right]_{\mathbf{c}^{-i}}=o(\epsilon_{\mathbf{x}})\to 0\quad\text{as }\epsilon_{\mathbf{x}}\to 0.$ (57)
Together with Eq. (49), we get:
$\mathbf{f}_{x}(\mathbf{x}_{1})-\mathbf{f}_{x}(\mathbf{x}_{0})\approx\mathbf{A}\mathbf{B}^{i}\left(\left[\log p(c^{i}\mid\mathbf{x}_{1})\right]_{c^{i}}-\left[\log p(c^{i}\mid\mathbf{x}_{0})\right]_{c^{i}}\right).$ (58)
@@
Again, when $\epsilon_{\mathbf{x}}\to 0$, i.e., when $p(\mathbf{c}\mid\mathbf{x})$ becomes sharply peaked at $\mathbf{c}_{\mathbf{x}}^{*}$, we can approximate:
$p(\mathbf{c}\mid\mathbf{x})\approx p(c^{i}\mid\mathbf{x})\cdot p(\mathbf{c}^{-i}\mid\mathbf{x}),$ (59)
Given the above, neglecting the constant term in the result in Theorem 3.1,
$\mathbf{f}_{x}(\mathbf{x})\approx\mathbf{A}\left[\log p(\mathbf{c}=\mathbf{c}_{i}\mid\mathbf{x})\right]_{\mathbf{c}_{i}},$ (60)
can be rewritten as
$\mathbf{f}_{x}(\mathbf{x})\approx\mathbf{A}\left[\log p(\mathbf{c}=\mathbf{c}_{i}\mid\mathbf{x})\right]_{\mathbf{c}_{i}}\approx\mathbf{A}(\mathbf{B}^{i}\left[\log p(c^{i}\mid\mathbf{x})\right]_{c^{i}}+\mathbf{B}^{-i}\left[\log p(\mathbf{c}^{-i}\mid\mathbf{x})\right]_{\mathbf{c}^{-i}}).$ (61)
In this case, for a data pair $(\mathbf{x}_{0},\mathbf{x}_{1})$ that differ only in the latent variable ${c}^{i}$, the representations $\mathbf{f}_{\mathbf{x}}(\mathbf{x})$ are passed to a linear classifier with weights $\mathbf{W}$.
@@
$\textbf{logits}\approx\mathbf{W}\big{(}\mathbf{A}(\mathbf{B}^{i}\left[\log p(c^{i}\mid\mathbf{x})\right]_{c^{i}}+\mathbf{B}^{-i}\left[\log p(\mathbf{c}^{-i}\mid\mathbf{x})\right]_{\mathbf{c}^{-i}})\big{)}.$ (62)
For correct classification under cross entropy loss, the logits must match the true probabilities:
$\textbf{logits}=s\left[p(c^{i}\mid\mathbf{x})\right]_{c^{i}},$ (63)
where $s$ is a scaling, corresponding to normalization before applying softmax in cross entropy loss.
@@
$\mathbf{W}(\mathbf{A}\mathbf{B}^{i})\approx s\mathbf{I},$ (64)